Q1: Given xi,T2, , and sample second moment is2. Compute the sample variance ,Xn. Suppose n = 10, sample first moment (i.e.,compute a 95% upper confidence interval of σ^2 .
Answer #1

The sample variance, s^2 = E(X^2) - E^2(X) = \frac{\sum_{i = 1}^{n} x_i^2}{n} - \overline{x}^2 = 2 - 1 = 1.

We know, \frac{(n-1)s^2}{\sigma^2} ~ \chi^2_{n-1}

The 95% upper confidence interval of \sigma^2 is = [0, \frac{(n-1)s^2}{\chi^2_{0.025}} ], where, n = 10, s^2 = 1

= [0, 3.33]. (Ans).